OpenStudy (anonymous):

\[7\sqrt{27/50}-3\sqrt{2/3}\]

OpenStudy (lgbasallote):

where are you stuck?

OpenStudy (anonymous):

I'm really not sure where to begin with this. Do I start by reducing the values in the radicals?

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

Is there even a common value between 27 and 50? Should I reduce them as their own radical, i.e. \[\sqrt{27} \sqrt{50}\]?

OpenStudy (lgbasallote):

find a factor

OpenStudy (lgbasallote):

factor out 27 and factor out 50

OpenStudy (lgbasallote):

can you think of a perfect square that fits in 27?

OpenStudy (anonymous):

3 cubed

OpenStudy (lgbasallote):

find a perfect square that you can divide 27 with...not cube

OpenStudy (anonymous):

I'm not sure I understand how 3 or 9 goes into 50

OpenStudy (lgbasallote):

stop thinking about 50

OpenStudy (anonymous):

am I treating 27/50 as a fraction?

OpenStudy (lgbasallote):

you're not going to turn this into simple terms

OpenStudy (lgbasallote):

the plan is to extract the roots of 27 and 50 SEPARATELY

OpenStudy (lgbasallote):

so \[\sqrt{27} = \sqrt{9 \times 3}\] what is sqrt of 9?

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

right. so \(\sqrt{27} \implies 3\sqrt 3\) agree?

OpenStudy (anonymous):

Should I reduce them as their own radical, i.e. 27−−√50−−√

OpenStudy (anonymous):

thats what I asked before and I was told no

OpenStudy (lgbasallote):

who said no?

OpenStudy (anonymous):

\[3\sqrt{3} and 5\sqrt{2}\]

OpenStudy (lgbasallote):

right so... \[\frac{\sqrt{27}}{\sqrt{50}} \implies \frac{3\sqrt 3}{5\sqrt 2}\]

OpenStudy (anonymous):

whoever the person was that bowed out from helping, 4th post I asked and porsha9 said no, so anyway, back on track

OpenStudy (lgbasallote):

anyways...you got what i just did?

OpenStudy (anonymous):

yes, thanks. so now do I multiply that by its conjugate? to remove the radical in the denominator

OpenStudy (lgbasallote):

no... multiply it to 7...because you originally had \[7 \sqrt{\frac{27}{50}}\] we just solved sqrt (27/50)

OpenStudy (anonymous):

so... \[\frac{ 21\sqrt{3} }{ 35\sqrt{2} }\] ?

OpenStudy (lgbasallote):

no...you don't multiply 7 to the denominator

OpenStudy (lgbasallote):

just the numerator

OpenStudy (anonymous):

Thanks for the help, it's getting late and I'm just going to stop at the learning assistance center tomorrow to work with someone on it. Thanks again.

OpenStudy (lgbasallote):

welcome